OpenStudy (anonymous):

A man divided $9,000 among his wife, son, and daughter. The wife received twice as much as the daughter, and the son received $1,000 more than the daughter. How much did each receive? If x is the amount the wife received, then which of the following expressions represents the amount received by the son?

OpenStudy (texaschic101):

x = 2d ----> x/2 = d s = d + 1000 s = x/2 + 1000 do you have answer choices ?

OpenStudy (anonymous):

yeah and that was correct thanks :)

OpenStudy (texaschic101):

good to hear....I wasn't 100% positive. Glad I could help